# Progressive Multi-Source Domain Adaptation for Personalized Facial Expression Recognition

### v1

@@ IX-C Visualization of Density-based Selection of Relevant Samples
Fig. 4 illustrates a t-SNE [69] representation of the embeddings from the last layer of the feature extractor ${F}(.)$ on Biovid target Subject-1 and Subject-5. We show the latent space of the features from a density-based selection of the replay domain before and after selecting a pertinent sample from a new source subject while adapting to a target subject. In the initial training step, the model retains more samples from the newly added source, as the network is more flexible in the initial stage, allowing more relevant samples to be included in a replay dictionary. On the other hand, as training progresses to step n, the replay dictionary is less affected by the newly added source subject. Two notable observations can be made here. First, the model gradually acquires useful knowledge from earlier subjects that are more closely related to the target, resulting in fewer samples being drawn from later sources, thereby minimizing the influence of more distant subjects. Second, the samples within the replay dictionary become tightly clustered with the target features, improving alignment with the target domain.

### v2

@@ X-E Visualization of Density-based Selection of Relevant Samples
Fig. 10 illustrates a t-SNE [72] representation of the embeddings from the last layer of the feature extractor F(.){F}(.) on Biovid target Subject-1 and Subject-5. We show the latent space of the features from a density-based selection of the replay domain before and after selecting a pertinent sample from a new source subject while adapting to a target subject. In the initial training step, the model retains more samples from the newly added source, as the network is more flexible in the initial stage, allowing more relevant samples to be included in a replay dictionary. On the other hand, as training progresses to step n, the replay dictionary is less affected by the newly added source subject. Two notable observations can be made here. First, the model gradually acquires useful knowledge from earlier subjects that are more closely related to the target, resulting in fewer samples being drawn from later sources, thereby minimizing the influence of more distant subjects. Second, the samples within the replay dictionary become tightly clustered with the target features, improving alignment with the target.
X-F Limitation and Failure Cases of P-MSDA
In addition to the cross-dataset results discussed in the main paper, we also analyzed failure cases in the in-the-wild datasets, Aff-Wild2 and BAH, where the absence of demographically aligned source subjects and highly imbalanced target distributions created significant challenges for P-MSDA. Table X reports the class distributions for all 10 target subjects of Aff-Wild2, where two types of failure cases emerge. Sub-9 highlights the effect of severe class imbalance, with the Fear class dominating (5,546 samples,  75% of the total), while Neutral (918), Anger (206), and Surprise (414) are relatively underrepresented, and both Disgust and Sadness are absent. This extreme skew biases pseudo-labeling and replay toward the majority class, limiting adaptation for minority expressions and contributing to Sub-9 reduced accuracy. In contrast, Sub-1 and Sub-3 represent failure cases caused not by imbalance but by large domain shifts, where no demographically or behaviorally similar source subjects exist. In these cases, the “closest” sources selected early in adaptation are poor matches, constraining transfer and resulting in weaker performance gains. These examples demonstrate that P-MSDA is sensitive to class imbalance (Sub-9) and to domain mismatch (Sub-1, Sub-3), challenges that are common in in-the-wild datasets such as Aff-Wild2. Similarly, in the BAH dataset, Sub-9 (a male East Asian subject aged 25–34) represents a critical failure case. As shown in Table XIII, the target subject belongs to a demographic group that is severely underrepresented in the source pool—only 3 East Asian males (25–34 years) are available compared to 26 White and 12 South Asian males of the same age group (Fig. 11). This imbalance constrained the similarity-based selection process, forcing adaptation to rely on less representative sources and producing only modest gains compared to other subjects. The demographic skew of BAH, where certain groups (e.g., White males) are heavily overrepresented while others are nearly absent, amplifies this limitation. These findings highlight that although P-MSDA demonstrates robustness across diverse conditions, its effectiveness diminishes when demographically relevant source subjects are missing, a scenario particularly common in real-world in-the-wild datasets.